Express each number as a product of its prime factors: 
140

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

140
By Taking the LCM of 140, we will get the product of its prime factor.
Therefore, 140 =2 × 2 × 5 × 7 × 1 = 2^2×5×7

Thanks